A horizontal pipe carries a smoothly flowing liquid of density of 1190 kg/m3. At Locations 1 and 2 along the pipe, the diameters are d1=6.63 cm and d2=2.45 cm, respectively.The flow speed at Location 1 is 2.33 m/s. What is the pressure difference ΔP between Location 2 and Location 1 (including its sign)? Ignore viscosity.
Answer #1

given Data = pipe is Hirizontal so (2,- Zq 8- 1190 kglm d, = 6.63x10²m da = 2,45 x16² m² blo o t using o 39 ag v= 2.33 milseeI (6.63 +1 5² x 2.33 = I (2.45x15 Xue ( 02217.06 milsee 17.66 -2-33 Ap= 169998